APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday August 18, 2010 (MJD 55427) ---=== OBSERVING TEAM ===--- Audrey Simmons Kaike Pan Paul Harding Stephen Snedden (support) ---=== OBSERVING PLAN ===--- Guider, engCam tests, possible Marvel Science! ---=== OBSERVING SUMMARY ===--- The night was lost to the clouds, high humidity. We did not get any science data, but had a few problems (see log body for details) 1. Rotator humming sound (indicating vibrations?). 2. Setting rotator fiducial (or rotator fiducial corrections) changes rotator position reading on mcp by ~ -90 degree. The fiducial correction for rot is turned off. 3. We collteced two sets rotator fiducial data. Analses indicates that some fiducial crossings are assigned to wrong fiducialID. 4. "c" and "init" commands on mcpMenu on enclosure laptop could not clear axes. 5. platedbActor was running but disconnected, and we could not reconnect it (error: Can't reconnect until invalid transaction is rolled back). ---=== OBSERVING LOG ===--- Afternoon: ---------- During afternoon check out, we found that Az fiducial at ~107 did not read out at both directions. We tried a few times and confirmed that rotator position will be changed by ~-90 degree if fiducial correction is reforced. No other problem to report. 00:40Z started ventilatin. 02:00 Stopped ventilation due to rain. Night: ------ 02:30Z Clouds and high humidity. 03:00Z we took some rotator fiducial data to check if a reasonable rotator fiducial table can be created. During the process we found that there are humming sound. This indicates a rotator vibration? we first moved the rotator at the zenith. Then we moved altitude down to 70 degree, we still heard the humming sound, but it is not as loud as at 90 degree. The other problem we had is on rotator position. During the rotator fiducial table data collecting process, when we set fiducial (the first fiducial cross), the rotator position reading on mcpMenu changed by ~ -90 degree. After our data collecting process, we set fiducial at different rotator positions, mcpMenu reading was changed by about the same amount, ~-90 degree. We also found that telescope axes could not be cleared (init or c) from mcpMenu or take quit a few tries, while "axis init" from tcc clear all axes without any problem. Steph analyzed the rotator fiducial data and found that some fiducial crossings are assigned to wrong fiducialID. To verify the physical zero on the rotator is correct, we collected another data data by moving the rotator to -180 degree, and zeroing it there. Then collect another set of rotator fiducial data, an anaysis indicates that the orignal assignment of rotator physical zero is correct. 08:30Z, we found that platedbActor was disconnected (It was connected in the evening). a platedb ping returned an error: platedb f text="Failed to connect to plateDB: Can't reconnect until invalid transaction is rolled back" We checked platedbActor status and found that it was running, tried to connect it, got the same error. Then we tried to restarted it, it returned the following error message: sdss3@hub25m-p[1003]$ stageManager platedbActor restart status Stopping platedbActor. /home/sdss3/products/Linux64/actorcore/v1_12_3/bin/stageManager: line 161: kill: (16272) - Operation not permitted NOT starting new platedbActor. Use restart if you want a new one. platedbActor is running as process 16272 We are not sure what kind of invalid transaction has been performed on the actor, and who did. We tried our best to fix it without success. Because weather will not allow us to open tonight, we decided not to call anyone, but send emails to related experts. 11:20Z endNight finsihed ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky ---=== TELESCOPE STATUS ===--- Never open. No interlocks bypassed. Status at: 11:20Z Telescope stowed at: 30,121,00 mount Instrument mounted: 11 Cartridge Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 210 lbs, 17.8 psi SP2 185 lbs, 20.2 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 TPM: tpm_v3_5_0 ----STUI software STUI: 1.0b2 2010-07-21 alerts: trunk+svn114742 apo: trunk+svn114701 boss: trunk+svn115542M boss.daq: 1.2.14 boss.specMech: Offline gcamera: trunk+svn115601M guider: branch_dstn-fiber-centers+svn115631 hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: trunk+svn111739 sop: v1_1_3 sos: ** FAILED ** tcc: 2.13.0 -----Other software sdssProcedure: v3_06 idlspec2d (sos3) --- v5_4_12 plate-mapper3 ------ NOCVS:v5_0_7 ---=== MIRROR NUMBERS ===--- ---=== PROBLEMS IN DETAIL ===---